Suppose that the mean and standard deviation of the scores on a statistics exam are 78 and 6.11, respectively, and are approximately normally distributed. Calculate the proportion of scores above 74.

Question 10 options:

1)

0.1920

2)

0.8080

3)

0.7437

4)

We do not have enough information to calculate the value.

5)

0.2563
Answer #1

Solution :

Given that ,

mean = = 78

standard deviation = = 6.11

P(x > 74) = 1 - p( x< 74)

=1- p P[(x - ) / < (74 - 78) / 6.11]

=1- P(z < -0.6547)

Using z table,

= 1 - 0.2563

= 0.7437

correct option is = 3